Q1: Find Vc(t) fort> 0 using general solution. + 10V 20V 0.15F 5Ω
Answer #1

to ov + lov Yello) sur 0.157 toy We under txo switch is at position ② and at inis posption erruit is in steady state state caNon Velo) 20i ei 32-0 2X10 7 -20 7 Vafo =171428 V all of S Cabautor Cannot change instantly ve lon) = vel0+) Evelo) : 17.1428aupply Kut in the loop 10t sila) + (a) = 0 10 film) - (a) Voltage a cenoss Capacitor. Velm) - (2n) 1.foo) =? : 2 x 10 v Ven)to 0.15 fo 7 1.5 T see 7 t/e Vill) Velo) t eneration solution of Caparitor rootege + (veros --Velas) 2.8571 + (17.1426 - 2.85